It took Jan y hours to drive 200 km. If she had increased her speed by 10km/h and driven for 2 h less, how far could she have gone? (Hint: Make a chart. Answer in terms of y.)
Answer by mananth(16949)

It took Jan y hours to drive 200 km. If she had increased her speed by 10km/h and driven for 2 h less, how far could she have gone? (Hint: Make a chart. Answer in terms of y.)
..
d= 200
time = y hours
...
speed = 200/y
...
increase speed by 10
so speed = (200/y)+10
..
traveled 2 hours less than before
y-2 hours
..
distance = speed * time
d= ((200/y)+10)*(y-2) km
...
simplified
d=(200+10y)/y * (y-2)
d=(200+10y)(y-2)/y
d=10(20+y)(y-2)/y km
...
m.ananth@hotmail.ca